OpenStudy (anonymous):

Find the standard form of the equation of the line that passes through the points (-1,-1) and (3,4)

OpenStudy (anonymous):

Did you find the slope of the line?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you know the formula for the slope?

OpenStudy (anonymous):

x1-y2 X2 -Y2 ?

OpenStudy (anonymous):

y1

OpenStudy (anonymous):

Not quite. \[Slope =\frac{\text{Change in y}}{\text{Change in x}} = \frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

so the slope is 3/5?

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

\[\frac{4-(-1)}{3-(-1)}\]

OpenStudy (anonymous):

so then what is my standard form equation?

OpenStudy (anonymous):

Once you have the slope, you can plug that into the point slope formula. \[y-y_1 = M(x-x_1)\] Where M is the slope.

OpenStudy (anonymous):

And \(x_1,y_1\) are the x and y values of a point on the line (you can pick either point for this)

OpenStudy (anonymous):

i give up

OpenStudy (anonymous):

Why? You were doing fine so far.

OpenStudy (anonymous):

You found M = 5/4

OpenStudy (anonymous):

You are given the x and y values for 2 points, just pick one and plug them in.